) Perform the indicated operations and write the answer in scientific notation:
(5.2 x 10^3)(4.5 x 10^5)

Found 2 solutions by Alan3354, josmiceli:
Answer by Alan3354(69443)

) Perform the indicated operations and write the answer in scientific notation:
(5.2 x 10^3)(4.5 x 10^5)
=====================
Use * (Shift 8) for multiply, not x.
(5.2*10^3)(4.5*10^5)
= 5.2*4.5*10^8
= 23.4*10^8
= 2.34*10^9
=========================
e format is commonly used, and saves some typing.
(5.2*10^3)(4.5*10^5)
(5.2e3)(4.5e5)
= 23.4e8
= 2.34e9

Answer by josmiceli(19441)